# Beyond Cosine Decay: On the effectiveness of Infinite Learning Rate Schedule for Continual Pre-training

### v2

@@ Abstract
The ever-growing availability of unlabeled data presents both opportunities and challenges for training artificial intelligence systems. While self-supervised learning (SSL) has emerged as a powerful paradigm for extracting meaningful representations from vast amounts of unlabeled data, existing methods still struggle to adapt to the non-stationary, non-IID nature of real-world data streams without forgetting previously learned knowledge. Recent works have adopted a repeated cosine annealing schedule for large-scale continual pre-training; however, these schedules (1) inherently cause forgetting during the re-warming phase and (2) have not been systematically compared to existing continual SSL methods. In this work, we systematically compare the widely used cosine schedule with the recently proposed infinite learning rate schedule and empirically find the latter to be a more effective alternative. Our extensive empirical evaluation across diverse image and language datasets demonstrates that the infinite learning rate schedule consistently enhances continual pre-training performance compared to a repeated cosine decay without being restricted to a fixed iteration budget. For instance, in a small-scale MAE pre-training setup, it outperforms several strong baselines from the literature. We then scale up our experiments to larger MAE pre-training and autoregressive language model pre-training. Our results show that the infinite learning rate schedule remains effective at scale, surpassing repeated cosine decay for both MAE pre-training and zero-shot LM benchmarks.
@@ 1 Introduction
Most approaches for continually pre-training foundation models typically utilize a repeated cosine annealing schedule (Loshchilov & Hutter, 2017) with fixed duration (Gupta et al., 2023; Defazio et al., 2023; Ibrahim et al., 2024; Parmar et al., 2024; Guo et al., 2024). Firstly, this implicitly assumes a terminal point in the training process, which severely limits the future pre-training on new datasets without undergoing significant forgetting. This fundamental limitation inhibits true continuous adaptation, as traditional learning rate schedules inevitably decay to near-zero values, effectively preventing further meaningful updates to the model. Secondly, re-warming the learning rate from its minimum value causes instability and exacerbates forgetting (Ibrahim et al., 2024). To overcome this constraint, recent works have explored more flexible infinite learning rate schedules that accommodate varying training durations (Zhai et al., 2022b; Defazio et al., 2024; Hu et al., 2024; Shen et al., 2024; Hägele et al., 2024). While these innovations emerged primarily from data-scaling research, their applications have begun to extend into CL, as demonstrated in (Garg et al., 2024; Ibrahim et al., 2024).
@@
In this work, we answer this question by comprehensively analyzingg the importance of learning rate schedules for self-supervised continual pre-training. Through extensive experiments across vision and language modalities, we believe that, to the best of our knowledge, we are the first to conduct a detailed comparison of infinite learning rate schedules with repeated annealing. Our results show that infinite schedules provides effective control over catastrophic forgetting in the non-IID setting, maintaining model performance across diverse data distributions. This work makes several key contributions:
@@
  • We demonstrate that infinite learning rate schedules alone and combined with experience replay outperform a number of sophisticated continual baselines in the context of self-supervised continual learning
  • We further demonstrate that, across multiple sequential large-scale vision and language pre-training tasks, infinite learning rate schedules outperform repeated cosine annealing.
  • Our results show that the Infinite Cosine Schedule should be the de facto schedule for continually pre-training foundation models due to its improved knowledge retention, relative to repeated cosine, in challenging non-IID self-supervised learning scenarios across both vision and language models.
@@ 2 Related Work
Continual pre-training (CPT) of Vision Foundation Models Continually pre-training Vision Transformers (ViTs) (Dosovitskiy et al., 2020; Bao et al., 2021) aims to adapt them to sequential data while mitigating catastrophic forgetting. Wang et al. (2022a) introduced the Lifelong Vision Transformer (LVT), incorporating inter-task attention to preserve critical weights across tasks. Ye & Bors (2024) proposed a task-free dynamic sparse ViT for scenarios without explicit task boundaries. The rise of large-scale foundation models has reshaped CL, particularly Vision-Language Models (VLMs) (Radford et al., 2021; Garg et al., 2024; Zhang et al., 2024; Singh et al., 2024), where CPT offers an efficient alternative to full retraining. Unlike parameter-efficient adaptation methods (Wang et al., 2022c; b; Smith et al., 2023), our work focuses on adapting the whole model.
@@ 3 The Need for Infinite Learning Rate Scheduling: Why It Matters?
Cosine Scheduling has been the de facto learning rate scheduling method for training large-scale models. Within a single cycle, it effectively balances between utilizing high learning rates for rapid optimization and gradually reducing rates to stabilize convergence, with the cycle duration expanding proportionally to the number of training iterations. However, this approach requires knowing the specific number of iterations in advance, which inhibits the ability to continually train a model on new incoming data. Several works (Zhai et al., 2022a; Hägele et al., 2024; Ibrahim et al., 2024; Hu et al., 2024) suggest alternatives such as infinite learning rate schedules and warmup-stable-decay(WSD), which offer the flexibility of training without predetermined step counts. These methods inherently support CL by default. An additional advantage of these alternate schedules is the ability to anneal or rapidly decrease the learning rate near the end of the training phase. This steep reduction has been shown to dramatically decrease the loss (Schaipp et al., 2025), enabling practitioners to preserve model checkpoints just before the decay phase for subsequent continual training cycles. This approach is particularly valuable in scenarios where the best-performing model must be deployed to users while anticipating the acquisition of additional high-quality data in the future.
@@
Inf $\begin{array}[]{c}\text{Inf Cosine}(n)=\begin{cases}\frac{n}{N_{w}}\cdot\eta_{max},&\text{if }n<N_{w}\\[10.0pt] \eta_{\textit{const}}+\frac{\eta_{\textit{max}}-\eta_{\textit{const}}}{2}\cdot\left(1+\cos\left(\pi\frac{n-N_{w}}{N_{c}-N_{w}}\right)\right),&\text{if }N_{\text{w}}<n\leq N_{\text{c}}\\[10.0pt] \eta_{const},&\text{if }N_{\text{c}}<n\leq N_{\text{d}}\\[10.0pt] \eta_{\textit{const}}\cdot\left(\frac{\eta_{\textit{min}}}{\eta_{\textit{const}}}\right)^{\frac{n-N_{d}}{t_{a}+N_{d}}}&\text{if }n>N_{d}\end{cases}\end{array}$ (1)
where $n$ is the current training step, $\eta_{max}$ and $\eta_{min}$ denote the maximum and minimum learning rates respectively, and $N_{w},N_{c}$ $N_{d}$ denote number of warmup steps, cooldown steps, and decay steps respectively, each specifying the transition points between the phases. $t_{a}$ denotes the amount of annealing steps required to achieve a converged checkpoint. As illustrated in Figure 1, the infinite learning rate schedule progresses through four distinct phases: during the warmup phase (until $N_{w}$), the learning rate increases linearly from 00 to $\eta_{max}$. It then transitions to a cosine cooldown schedule until reaching $N_{c}$, followed by a stable constant phase until $N_{d}$, before finally annealing exponentially until the end of the task. We additionally define the cooldown proportion as $P=N_{c}/N$. For continually pre-training on subsequent tasks, we can efficiently resume training from the pre-annealed checkpoint at the constant learning rate $\eta_{const}$. Note that the subsequent tasks will only consist of constant phase and annealing, eliminating the need for rewarming.
@@ 4.1 Continual pre-training of MAEs
Baselines and Adaptations: We compare our approach with the following CL baselines, adapting them for self-supervised pre-training: Sequential Fine-tuning: trains sequentially without mitigating forgetting, serving as the primary baseline. Experience Replay (ER) (Rolnick et al., 2019): maintains a memory buffer with {40%, 50%} samples of prior tasks, sampled uniformly. Each batch contains equal proportion of current task data and randomly sampled data from replay buffer. Memory Aware Synapses (MAS) (Aljundi et al., 2018): adapted for self-supervised learning by computing importance of weights from the L2 norm of the encoder’s output, with a regularization $\lambda=0.75$. Learning without Forgetting (LwF) (Li & Hoiem, 2017): modified for self-supervised learning with feature distillation on the encoder’s output, weighted by $\alpha=0.75$. GDumb (Prabhu et al., 2020): Uses stratified sampling to maintain a balanced buffer. The model resets to random initialization for each new task and trains from scratch on buffer data. For evaluation, we use standard CL metrics from Lopez-Paz & Ranzato (2017): Average Accuracy (Acc), Forward Transfer (FWT), and Backward Transfer (BWT), defined in Appendix A.3.
@@ 4.3 Continually pre-training LLMs
All the three datasets are tokenized through LLaMA-3 tokenizer (Grattafiori et al., 2024) owing to its large vocabulary size of 128K tokens (100K from titktoken [^2] https://github.com/openai/tiktoken/tree/main and 28K additional tokens for non-English languages). We sample a small subset of 100B tokens from each of the DCLM-Baseline (total = 3T), Stack (total = 744B), and OSCAR (total = 168B) datasets for our CPT experiments. We would like to emphasize that as the domain shifts farther away from the tokenizer’s training corpus, the tokenizer might become the key bottleneck to performance. Such scenarios would be unrealistic without a way to adapt the tokenizer. With this in mind, we were careful to select challenging new domains that are still well represented in the tokenizer’s vocabulary. We leave the treatment continual tokenizer adaptation to future work.
@@ 5.1 Results for pre-training MAE on CIFAR10
Table 1: Average linear probe accuracy (Acc) and Backward Transfer (BWT) (where $\uparrow$ indicates that higher is better) for comparing CL baselines utilizing cosine schedule with Infinite Schedule on CIFAR10 with varying replay (ER) strategies. It can be observed that the infinite schedule (Inf Cos) consistently achieves superior performance compared to the cosine
@@
Interestingly, in this setup, the combination of the repeated cosine schedule and experience replay (ER) degrades model performance, as seen in the comparison between FT-seq and ER with 40% replay. This decline likely stems from limited data diversity in small datasets, leading the more aggressive re-warming of the repeated cosine schedule to overfitting to the replay buffer. In contrast, the infinite learning rate schedule eliminates rewarming, effectively circumventing these issues.While replay behavior in small data scenarios is not our primary focus, it is worth noting that we used a relatively large replay buffer despite the dataset’s limited size. The key finding is that the infinite cosine schedule, despite its simplicity, consistently outperforms baselines in both average accuracy and backward transfer (BWT). Notably, the strong performance gains with larger replay buffers suggest that our method scales effectively to large-scale pre-training, where the vast size of modern datasets provides sufficient replay samples to mitigate catastrophic forgetting, even at low buffer sampling rates.
@@
Table 2: Performance comparison between cosine (Cos) and infinite cosine (Inf) for MAE pre-training across different tasks, with and without a replay buffer. Grey values indicate performance on datasets that were unseen during training at that stage. Each row shows model performance after the model has completed training on the task specified in the row label. The infinite schedule generally preserves knowledge better, particularly in the presence of multiple distribution shifts. Note that this is shown by the superior knowledge retention (bolded) on the previous tasks after learning new tasks. The table also presents key metrics: Average Accuracy (Avg. Acc.), Forward Transfer (FWT), and Backward Transfer (BWT), where $\uparrow$ indicates that higher is better.
@@ 5.2 Results for pre-training mae on multiple datasets
We present the results of our experiments on large scale MAE pre-training in Table 2 (left). The infinite schedule achieves accuracy comparable to a cosine schedule after ImageNet ($\mathcal{D}_{0}$) pre-training. The effectiveness becomes more pronounced after continual training on Places2 ($\mathcal{D}_{1}$) with a replay buffer (ER), where the infinite schedule outperforms the cosine schedule on the previous task while achieving better performance on the current dataset. Even under the strong distribution shift introduced by Firerisk ($\mathcal{D}_{2}$), the infinite cosine schedule proves remarkably robust, achieving 57.50% accuracy on ImageNet. After completing all three tasks, the infinite schedule achieves an average accuracy of 50.18% across all datasets, $\approx$ 1.3% higher than the cosine schedule. The Forward Transfer (FWT) metrics are comparable between the two schedules, while the infinite schedule shows better resistance to catastrophic forgetting with a higher Backward Transfer (BWT).
@@
Effect of cooldown proportion ($P$) and constant learning rate ($\eta_{const}$): In Figure 2, we analyze how the cooldown proportion and constant learning rate in the infinite schedule affect model performance on past and current tasks in ImageNet and Places2. The graphs compare linear probe validation loss across epochs for the standard cosine schedule and infinite schedules with varying configurations. Our analysis shows that lower constant learning rate ($\eta_{const}=3.75e-5$) consistently reduces forgetting as compared to higher rate ($\eta_{const}=1.12e-4$), Further, it can be observed that for $\eta_{const}=3.75e-5$ cooldown proportion has negligble effect, but for $\eta_{const}=1.12e-4$ shorter cooldown period ($P=0.3$) outperform longer period ($P=0.5$). This is likely because shorter cooldown phase represents quick decay to a stable $\eta_{const}$ whereas a longer cooldown would mean a high learning rate for longer durations, which could cause instability in training, thus increasing forgetting. While the cosine schedule performs better on current tasks, this gap narrows with an appropriately low constant learning rate in the infinite schedule. A similar trend is observed in training without a replay buffer, as shown in Appendix B. nResults for Continual Pre-training LLMs

### v3

@@ Abstract
The ever-growing availability of unlabeled data presents both opportunities and challenges for training artificial intelligence systems. While self-supervised learning (SSL) has emerged as a powerful paradigm for extracting meaningful representations from vast amounts of unlabeled data, existing methods still struggle to adapt to the non-stationary, non-IID nature of real-world data streams without forgetting previously learned knowledge. Recent works have adopted a repeated cosine annealing schedule for large-scale continual pre-training; however, these schedules (1) inherently cause forgetting during the re-warming phase and (2) have not been systematically compared to existing continual SSL methods. In this work, we systematically compare the widely used cosine schedule with the recently proposed infinite learning rate schedule and empirically find the latter to be a more effective alternative. Our extensive empirical evaluation across diverse image and language datasets demonstrates that the infinite learning rate schedule serves as a flexible and robust alternative without being restricted to a fixed iteration budget. For instance, in a small-scale Masked AutoEncoder (MAE) pre-training setup, it performs competitively against several strong baselines from the literature. We then scale up our experiments to larger MAE pre-training and autoregressive language model pre-training. Our results show that the infinite learning rate schedule continues to show stable and effective behavior—maintaining knowledge retention and adaptability across tasks, while eliminating the need to know the total training iterations in order to fix a termination point. Code is available at https://github.com/Pauljanson002/beyond-cosine.
@@ 1 Introduction
Most approaches for continually pre-training foundation models typically utilize a repeated cosine annealing schedule (Loshchilov & Hutter, 2017) with fixed duration (Gupta et al., 2023; Defazio et al., 2023; Ibrahim et al., 2024; Parmar et al., 2024; Guo et al., 2024). In this schedule, the learning rate undergoes an initial warmup phase to reach its maximum value, followed by a cosine decay that gradually reduces it to the minimum value precisely at the end of training (purple in Figure 1). This implicitly assumes a terminal point in the training process, which severely limits the future pre-training on new datasets without undergoing significant forgetting. This fundamental limitation inhibits true continuous adaptation, as traditional learning rate schedules inevitably decay to near-zero values, effectively preventing further meaningful updates to the model. Secondly, re-warming the learning rate from its minimum value causes instability and exacerbates forgetting (Ibrahim et al., 2024). To overcome this constraint, recent works have explored more flexible infinite learning rate schedules that accommodate varying training durations (Zhai et al., 2022b; Defazio et al., 2024; Hu et al., 2024; Shen et al., 2024; Hägele et al., 2024). These schedules consist of four distinct phases: an initial warmup, followed by a decay phase (e.g., cosine, inverse square root) that reduces the learning rate to a constant value, a plateau phase where this rate is maintained, and a final rapid annealing phase near the end of training (black line in Figure 1). While these innovations emerged primarily from data-scaling research, their applications have begun to extend into CL, as demonstrated in (Garg et al., 2024; Ibrahim et al., 2024).
@@
In this work, we answer this question by comprehensively analyzing the importance of learning rate schedules for self-supervised continual pre-training. Through extensive experiments across vision and language modalities, we believe that, to the best of our knowledge, we are the first to conduct a detailed comparison of infinite learning rate schedules with repeated annealing. Our results show that infinite schedules offer a flexible and stable approach to mitigating catastrophic forgetting in the non-IID setting, maintaining model performance across diverse data distributions, and offering a competitive advantage in retaining the previous task performance as compared to the repeated cosine with warmup.
This work makes several key contributions:
@@
  • We demonstrate that just using infinite learning rate schedules alone and combined with experience replay can outperform a number of sophisticated continual baselines in the context of self-supervised continual learning (Sec 5.1)
  • We further demonstrate that, across multiple sequential large-scale vision and language pre-training tasks, infinite learning rate schedules either match or outperform repeated cosine annealing, while offering greater flexibility by not requiring a predefined number of data(Sec 5.2, 5.3).
  • Our results show that the Infinite Cosine Schedule is a flexible and robust alternative to repeated cosine decay for continually pre-training foundation models, demonstrating that it can improve knowledge retention and adaptability relative to the repeated cosine schedule, in challenging non-IID self-supervised learning scenarios across both vision and language modalities.
@@ 2 Related Work
Continual pre-training (CPT) of Vision Foundation Models Continually pre-training Vision Transformers (ViTs) (Dosovitskiy et al., 2020; Bao et al., 2021) aims to adapt them to sequential data while mitigating catastrophic forgetting. Wang et al. (2022a) introduced the Lifelong Vision Transformer (LVT), incorporating inter-task attention to preserve critical weights across tasks. Ye & Bors (2024) proposed a task-free dynamic sparse ViT for scenarios without explicit task boundaries. The rise of large-scale foundation models has reshaped CL, particularly Vision-Language Models (VLMs) (Radford et al., 2021; Janson et al., 2022; Garg et al., 2024; Zhang et al., 2024; Singh et al., 2024), where CPT offers an efficient alternative to full retraining. Unlike parameter-efficient adaptation methods (Wang et al., 2022c; b; Smith et al., 2023), our work focuses on adapting the whole model.
@@ 3 The Need for Infinite Learning Rate Scheduling: Why It Matters?
Cosine Scheduling has been the de facto learning rate scheduling method for training large-scale models. Within a single cycle, it effectively balances between utilizing high learning rates for rapid optimization and gradually reducing rates to stabilize convergence, with the cycle duration expanding proportionally to the number of training iterations. However, this approach requires knowing the specific number of iterations in advance, which inhibits the ability to continually train a model on new incoming data. Several works (Zhai et al., 2022a; Hägele et al., 2024; Ibrahim et al., 2024; Hu et al., 2024) suggest alternatives such as infinite learning rate schedules and warmup-stable-decay(WSD), which offer the flexibility of training without predetermined step counts. These methods inherently support CL by default. An additional advantage of these alternate schedules is the ability to anneal or rapidly decrease the learning rate near the end of the training phase. This steep reduction has been shown to dramatically decrease the loss (Schaipp et al., 2025), enabling practitioners to preserve model checkpoints just before the decay phase for subsequent continual training cycles. Further infinite schedules enable targeted replay strategies, where replay buffer configurations can be adjusted dynamically (e.g., during annealing) to emphasize retention of past knowledge, providing more control than fixed-schedule approaches. This approach is particularly valuable in scenarios where the best-performing model must be deployed to users while anticipating the acquisition of additional high-quality data in the future, as it enables continued training without a fixed iteration budget, supports agile checkpointing for seamless updates, and allows replay strategies to be tuned dynamically for improved knowledge retention across tasks.
@@
$\begin{array}[]{c}\text{Inf Cosine}(n)=\begin{cases}\frac{n}{N_{w}}\cdot\eta_{max},&\text{if }n<N_{w}\\[10.0pt] \eta_{\textit{const}}+\frac{\eta_{\textit{max}}-\eta_{\textit{const}}}{2}\cdot\left(1+\cos\left(\pi\frac{n-N_{w}}{N_{c}-N_{w}}\right)\right),&\text{if }N_{\text{w}}<n\leq N_{\text{c}}\\[10.0pt] \eta_{const},&\text{if }N_{\text{c}}<n\leq N_{\text{d}}\\[10.0pt] \eta_{\textit{const}}\cdot\left(\frac{\eta_{\textit{min}}}{\eta_{\textit{const}}}\right)^{\frac{n-N_{d}}{t_{a}+N_{d}}}&\text{if }n>N_{d}\end{cases}\end{array}$ (1)
where $n$ is the current training step, $\eta_{max}$ and $\eta_{min}$ denote the maximum and minimum learning rates respectively, and $N_{w},N_{c}$ $N_{d}$ denote number of warmup steps, cooldown steps, and decay steps respectively, each specifying the transition points between the phases. $t_{a}$ denotes the amount of annealing steps required to achieve a converged checkpoint. As illustrated in Figure 1, the infinite learning rate schedule progresses through four distinct phases: during the warmup phase (until $N_{w}$), the learning rate increases linearly from 0 to $\eta_{max}$. It then transitions to a cosine cooldown schedule until reaching $N_{c}$, followed by a stable constant phase until $N_{d}$, before finally annealing exponentially until the end of the task. We additionally define the cooldown proportion as $P=N_{c}/N$. For continually pre-training on subsequent tasks, we can efficiently resume training from the pre-annealed checkpoint at the constant learning rate $\eta_{const}$. Note that the subsequent tasks will only consist of constant phase and annealing, eliminating the need for rewarming.
@@ 4.1 Continual pre-training of MAEs
Baselines and Adaptations: We compare our approach with the following CL baselines, adapting them for self-supervised pre-training: Sequential Fine-tuning: trains sequentially without mitigating forgetting, serving as the primary baseline. Experience Replay (ER) (Rolnick et al., 2019): maintains a memory buffer with {40%, 50%} samples of prior tasks, sampled uniformly. Each batch contains equal proportion of current task data and randomly sampled data from replay buffer. Memory Aware Synapses (MAS) (Aljundi et al., 2018): estimates parameter importance by measuring how changes affect the model output, then penalizes updates to important weights. We adapted it for self-supervised learning by computing importance of weights from the L2 norm of the encoder’s output, with a regularization $\lambda=0.75$. Learning without Forgetting (LwF) (Li & Hoiem, 2017): preserves knowledge by distilling responses from the previous model version. We modified it for self-supervised learning with feature distillation on the encoder’s output, weighted by $\alpha=0.75$. GDumb (Prabhu et al., 2020): Uses stratified sampling to maintain a balanced buffer. The model resets to random initialization for each new task and trains from scratch on buffer data. For evaluation, we use standard CL metrics from Lopez-Paz & Ranzato (2017): Average Accuracy (Acc), Forward Transfer (FWT), and Backward Transfer (BWT), defined in Appendix A.3.
@@ 4.3 Continually pre-training LLMs
All the three datasets are tokenized through LLaMA-3 tokenizer (Grattafiori et al., 2024) owing to its large vocabulary size of 128K tokens (100K from titktoken [^2] https://github.com/openai/tiktoken/tree/main and 28K additional tokens for non-English languages). We sample a small subset of 100B tokens from each of the DCLM-Baseline (total = 3T), Stack (total = 744B), and OSCAR (total = 168B) datasets for our CPT experiments. We would like to emphasize that as the domain shifts farther away from the tokenizer’s training corpus, the tokenizer might become the key bottleneck to performance. Such scenarios would be unrealistic without a way to adapt the tokenizer. With this in mind, we were careful to select challenging new domains that are still well represented in the tokenizer’s vocabulary. Though we did not perform a formal tokenizer coverage analysis, our use of German alongside English datasets aligns with LLaMA-3’s multilingual capabilities Grattafiori et al. (2024). Stable validation loss across domains indicates no significant tokenizer-data mismatch in practice. We leave the treatment, continual tokenizer adaptation to future work.
@@ 5.1 Results for pre-training MAE on CIFAR10
Table 1: Average linear probe accuracy (Acc) and Backward Transfer (BWT) (where $\uparrow$ indicates that higher is better) for comparing CL baselines utilizing cosine schedule with Infinite Schedule on CIFAR10 with varying replay (ER) strategies. It can be observed that the infinite schedule (Inf Cos) consistently achieves superior performance compared to the cosine schedule across all experimental configurations.
@@
Interestingly, in this setup, the combination of the repeated cosine schedule and experience replay (ER) degrades model performance, as seen in the comparison between FT-seq and ER with 40% replay. This decline likely stems from limited data diversity in small datasets, leading the more aggressive re-warming of the repeated cosine schedule to overfitting to the replay buffer. In contrast, the infinite learning rate schedule eliminates rewarming, effectively circumventing these issues. We would like to emphasize that the unexpectedly poor performance of these methods relative to FT-seq stems from a fundamental difference in the continual learning setting. While these baselines were originally designed for incremental supervised classification, our work centers on incremental masked image modeling (MIM) — a self-supervised task with distinct objectives and evaluation protocols. To the best of our knowledge, this is the first systematic evaluation of continual learning baselines in a self-supervised MIM setting, leaving us without alternatives specifically designed for this novel paradigm. Therefore, the results should be interpreted with this context in mind.
While replay behavior in small data scenarios is not our primary focus, it is worth noting that we used a relatively large replay buffer despite the dataset’s limited size. The key finding is that the infinite cosine schedule, despite its simplicity, consistently outperforms baselines in both average accuracy and backward transfer (BWT). Notably, the strong performance gains with larger replay buffers suggest that our method scales effectively to large-scale pre-training, where the vast size of modern datasets provides sufficient replay samples to mitigate catastrophic forgetting, even at low buffer sampling rates.
@@
Table 2: Performance comparison between cosine (Cos) and infinite cosine (Inf) for MAE pre-training across different tasks, with and without a replay buffer. Grey values indicate performance on datasets that were unseen during training at that stage. Each row shows model performance after the model has completed training on the task specified in the row label. The infinite schedule generally preserves knowledge better, particularly in the presence of multiple distribution shifts. Note that this is shown by the superior knowledge retention (bolded) on the previous tasks after learning new tasks. The bottom section presents key averaged metrics across all three tasks: Average Accuracy (Avg. Acc.), Forward Transfer (FWT), and Backward Transfer (BWT), (where $\uparrow$ indicates that higher is better). Infinite cosine achieves better overall results, especially in reducing forgetting (as shown by less negative BWT values).
@@ 5.2 Results for pre-training mae on multiple datasets
We present the results of our experiments on large scale MAE pre-training in Table 2 (left). The infinite schedule achieves accuracy comparable to a cosine schedule after ImageNet ($\mathcal{D}_{0}$) pre-training. The effectiveness becomes more pronounced after continual training on Places2 ($\mathcal{D}_{1}$) with a replay buffer (ER), where the infinite schedule outperforms the cosine schedule on the previous task while achieving better performance on the current dataset. Even under the strong distribution shift introduced by Firerisk ($\mathcal{D}_{2}$), the infinite cosine schedule proves remarkably robust, achieving 57.50% accuracy on ImageNet. After completing all three tasks, the infinite schedule achieves an average accuracy of 50.18% across all datasets, $\approx$ 1.3% higher than the cosine schedule. The Forward Transfer (FWT) metrics are comparable between the two schedules, while the infinite schedule shows better resistance to catastrophic forgetting with a higher Backward Transfer (BWT). We perform extended forgetting analysis in Appendix H.
@@
Effect of cooldown proportion ($P$) and constant learning rate ($\eta_{const}$): In Figure 2, we analyze how the cooldown proportion and constant learning rate in the infinite schedule affect model performance on past and current tasks in ImageNet and Places2. The graphs compare linear probe validation loss across epochs for the standard cosine schedule and infinite schedules with varying configurations. Our analysis shows that lower constant learning rate ($\eta_{const}=3.75e-5$) consistently reduces forgetting as compared to higher rate ($\eta_{const}=1.12e-4$), Further, it can be observed that for $\eta_{const}=3.75e-5$ cooldown proportion has negligble effect, but for $\eta_{const}=1.12e-4$ shorter cooldown period ($P=0.3$) outperform longer period ($P=0.5$). This is likely because shorter cooldown phase represents quick decay to a stable $\eta_{const}$ whereas a longer cooldown would mean a high learning rate for longer durations, which could cause instability in training, thus increasing forgetting. While the cosine schedule performs better on current tasks, this gap narrows with an appropriately low constant learning rate in the infinite schedule. A similar trend is observed in training without a replay buffer, as shown in Appendix B.
5.3 Results for Continual Pre-training LLMs